# Reconstructing Quantitative Cerebral Perfusion Images Directly From Measured Sinogram Data Acquired Using C-arm Cone-Beam CT

### v1

@@ Abstract
To shorten the door-to-puncture time for better treating patients with acute ischemic stroke, it is highly desired to obtain quantitative cerebral perfusion images using C-arm cone-beam computed tomography (CBCT) equipped in the interventional suite. However, limited by the slow gantry rotation speed, the temporal resolution and temporal sampling density of typical C-arm CBCT are much poorer than those of multi-detector-row CT in the diagnostic imaging suite. The current quantitative perfusion imaging includes two cascaded steps, time-resolved image reconstruction and perfusion parametric estimation. For time-resolved image reconstruction, technical challenge imposed by poor temporal resolution and poor sampling density causes inaccurate quantification of the temporal variation of cerebral artery and tissue attenuation values. For perfusion parametric estimation, it remains a technical challenge to appropriately design the handcrafted regularization for better solving the associated deconvolution problem. These two challenges together prevent from obtaining quantitatively accurate perfusion images using C-arm CBCT. The purpose of this work is to simultaneously address these two challenges via combining the two cascaded steps into a single joint optimization problem and reconstructing quantitative perfusion images directly from the measured sinogram data. In the developed direct cerebral perfusion parametric image reconstruction technique, TRAINER in short, the quantitative perfusion images have been represented as a subject-specific conditional generative model trained under the constraint of the time-resolved CT forward model, perfusion convolutional model, and the subject’s own measured sinogram data. Results shown in this paper demonstrated that using TRAINER, quantitative cerebral perfusion images can be accurately obtained using C-arm CBCT in the interventional suite.
@@ I-B Technical Challenges for Quantitative Cerebral Perfusion Imaging Using C-arm Cone-Beam CT
Several techniques have previously been proposed to address the potential challenges in time-resolved image reconstruction using C-arm CBCT. One scheme [7, 8, 9] was to decompose the time attenuation curves (TACs) of different image voxels in terms of known temporal basis functions such as Gaussian functions, gamma-variate functions, linear or cubic splines. The incorporation of known basis functions dramatically reduces the number of unknowns, hence, strongly regularizes the otherwise ill-posed optimization problem. In the second scheme, a temporal deconvolution method was proposed in [10] to improve the temporal resolution and temporal sampling density by deconvolving the reconstructed TACs with a known convolution kernel simulating the temporal resolution degradation scenario. In the third scheme, a local TACs optimization framework was proposed in [11] to model the TACs inside the vessel regions and optimize the TACs constrained by the acquired data. The fourth scheme was to improve temporal resolution via the limited-view image reconstruction strategy. One such method is the SMART-RECON [12] that reconstructs up to four time frames with temporal variation without suffering from severe limited-view artifacts using data acquired over the short-scan range [13]. By incorporating the prior knowledge of periodicity of the limited-view artifacts in the C-arm CBCT perfusion data acquisition scheme, eSMART-RECON [6] further improves the temporal resolution to 4-7.5 frames per second (fps). Fifth, by recasting the temporal resolution improvement task as a temporal extrapolation task, AIRPORT [14] improves the temporal resolution to 40 fps, the time window needed to acquire a single projection view data using typical C-arm CBCT and the data acquired over the short-scan range.
@@ I-D Technical Challenges for Perfusion Parametric Estimation
Perfusion parametric estimation aims to quantify cerebral blood flow (CBF), cerebral blood volume (CBV), mean transit time (MTT), and time to peak (TTP) etc using the reconstructed time-resolved CT images. For perfusion parametric estimation, currently available techniques [15, 16, 17, 18, 19, 20, 21] incorporate handcrafted regularization to heavily constrain the otherwise ill-posed deconvolution problem especially under nonideal data acquisition conditions such as low exposure levels. However, it remains a technical challenge to appropriately design the regularization for better solving the deconvolution problem. Inappropriate designs of regularizations may lead to bias in the obtained parametric images [22, 23] even if the used time-resolved images are quantitatively accurate.
@@ I-E Purpose and Innovations of This Work
Although significant efforts have been made to reconstruct time-resolved images and estimate perfusion parametric images from the reconstructed time-resolved CT images, there is no work published yet to address inaccurate perfusion imaging jointly caused by the poor temporal resolution of C-arm CBCT and current perfusion parametric estimation techniques due to biased regularization. These two challenges together prevent from obtaining quantitative accurate perfusion images using C-arm CBCT.
In the current quantitative perfusion imaging pipeline, time-resolved image reconstruction and perfusion parametric estimation are treated as two cascaded challenges. It is our conviction that addressing each of them independently may fail to achieve the needed overall accuracy. On the contrary, in this work, instead of addressing each challenge independently, we recast the quantitative perfusion imaging problem as a single joint optimization problem and reconstructing quantitative perfusion images directly from the measured sinogram data of each individual subject.
To implement this idea, we developed direct cerebral perfusion parametric image reconstruction technique, TRAINER in short, that jointly leverages the forward model of the time-resolved CT data acquisition and the convolutional model for cerebral perfusion parametric estimation. To avoid potentially biased regularization design for perfusion parametric estimation, the perfusion parametric images in TRAINER have been represented as a subject-specific conditional generative model trained under the constraint of subject’s own measured sinogram data.
@@
  • It explicitly leverages the measured sinogram data of each individual subject to guarantee the accuracy of perfusion parametric images for each individual subject;
@@
In this work, we demonstrated that using TRAINER, quantitative perfusion images can be accurately obtained via training a subject-specific conditional generative model under the constraint of the subject’s own measured sinogram data. Results shown in the paper demonstrated that the two technical challenges, i.e. poor temporal resolution of C-arm CBCT and inaccurate perfusion parametric estimation methods, have been simultaneously addressed using the proposed TRAINER technique. TRAINER may help quickly identify patients who are most likely benefit from EVT directly using C-arm CBCT in the interventional suite.
@@ II-A Forward Model for Time-Resolved CT Data Acquisition
Considering the ideal CT imaging scenario when the image object is static and stationary and a monochromatic X-ray spectrum is used in the data acquisition, the forward model under proper digitization can be formulated as the following linear system:
$\bar{\mathbf{y}}$ $=\mathbf{A}\mathbf{x},$ (1)
@@
$\bar{\mathbf{y}}_{t}$ $=\mathbf{A}_{t}\mathbf{x}_{t},$ (2)
where $\bar{\mathbf{y}}_{t}\in\mathbb{R}^{\frac{M}{K}\times 1}$, $\mathbf{x}_{t}\in\mathbb{R}^{N\times 1}$ denote the modeled line integrals values and the image to be reconstructed associated to the $t$-th time frame respectively, $\mathbf{A}_{t}\in\mathbb{R}^{\frac{M}{K}\times N}$ denotes CT system matrix corresponding to $t$-th time frame, $t=1,2,\cdots,T$ denotes the index of time frame. Here $T_{0}$ denotes the number of short-scan acquisitions, $K$ denotes the number of subsets within a single short-scan range, and $T=KT_{0}$ denotes the total number of time frames to be reconstructed.
Since our goal is to jointly reconstruct the dynamic image object $\{\mathbf{x}_{t}\}$, it is more convenient to stack up these linear sub-systems as follows:
@@
$\mathbf{X}$ $=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}],$ (5)
$\bar{\mathbf{Y}}$ $=[\bar{\mathbf{y}}_{1},\bar{\mathbf{y}}_{2},\cdots,\bar{\mathbf{y}}_{T}],$ (6)
where $\mathcal{A}\in\mathbb{R}^{MT_{0}\times NT}$ takes each system matrix $\mathbf{A}_{t}$ as its diagonal matrix block, $\mathbf{X}\in\mathbb{R}^{N\times T}$ denotes the time-resolved image matrix stacking each image associated with $t$-th time frame $\mathbf{x}_{t}$ along its column dimension, and $\vec{\cdot}$ denotes the vectorization operator.
@@ II-B Convolutional Model for Cerebral Perfusion Parametric Estimation
According to the convolutional model for cerebral perfusion parametric mapping [24, 25, 26], the relation between the time-resolved image matrix $\mathbf{X}$ and the flow-scaled residual matrix $\mathbf{C}$ can be represented as the following linear model:
$\mathbf{X}$ $=\mathbf{C}\mathbf{B}^{\text{T}}$ (7)
$B_{i,j}$ =$=\begin{cases}h_{\text{AIF}}(t_{i-j+1}),~{}~{}~{}j\leq i\cr 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}j>i\end{cases}$ (8)
$C_{i,j}$ =$=\begin{cases}\rho\cdot\text{CBF}_{i}\cdot\exp\left(-\left(t_{j}-\text{MTT}_{i}\right)\right),~{}t_{j}\geq\text{MTT}_{i}\cr 1,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}t_{j}<\text{MTT}_{i}\end{cases}$ (9)
where $\mathbf{B}\in\mathbb{R}^{T\times T}$, $B_{i,j}$ denotes the matrix element of $\mathbf{B}$, $h_{\text{AIF}}(t_{i})$ denotes the digitized arterial input function [27]. $C_{i,j}$ denotes the matrix element of $\mathbf{C}\in\mathbb{R}^{N\times T}$ which denotes the flow-scaled residual matrix stacking up each flow-scaled residual function values of all image voxels associated to the $j$-th time frame, $\mathbf{c}_{j}\in\mathbb{R}^{N\times 1}$, along its column direction. The matrix form $\mathbf{C}\mathbf{B}^{\text{T}}$ represents the digitized convolutional process of the arterial input function $h_{\text{AIF}}(t)$ and the flow-scaled residual function for each image voxel independently. $\rho$ denotes the nominal density of cerebral tissues. $\text{CBF}_{i}$ and $\text{MTT}_{i}$ denotes the $i$-th element of the two perfusion parametric images to be reconstructed, CBF and MTT, respectively.
@@ II-C Quantitative Cerebral Perfusion Image Reconstruction Directly From Measured Sinogram Data
The two perfusion parametric images to be reconstructed CBF and MTT can be obtained by solving the following optimization problem:
@@
$\mathbf{U}$ $:=\left[\mathbf{\text{CBF}},\mathbf{\text{MTT}}\right],$
where $\mathbf{W}\in\mathbb{R}^{MT_{0}\times MT_{0}}$ denotes the statistical weighting matrix and $\vec{\mathbf{Y}}\in\mathbb{R}^{MT_{0}\times 1}$ denotes the measured data stacking the measurements associated to each scan along its column direction.
@@ II-D Subject-Specific Conditional Generative Models for Perfusion Parametric Image Representation
To avoid potentially biased regularization design for perfusion parametric estimation, the perfusion parametric images in the proposed TRAINER have been represented as a subject-specific conditional generative model trained under the constraint of subject’s own measured sinogram data. We parameterize the unknowns (CBF and MTT) as follows:
@@
where $\mathcal{G}_{\mathbf{\Theta}}$ denote the parameterized perfusion images, CBF and MTT, from the noise image sampled from the uniform distribution, $\mathbf{Z}\in\mathbb{R}^{N\times 2}$. Inspired by the deep image prior [28], we represent $\mathcal{G}_{\mathbf{\Theta}}$ by typical convolutional neural networks, whose model parameters are denoted by $\mathbf{\Theta}$.

### v2

@@ Abstract
To shorten the door-to-puncture time for better treating patients with acute ischemic stroke, it is highly desired to obtain quantitative cerebral perfusion images using C-arm cone-beam computed tomography (CBCT) equipped in the interventional suite. However, limited by the slow gantry rotation speed, the temporal resolution and temporal sampling density of typical C-arm CBCT are much poorer than those of multi-detector-row CT in the diagnostic imaging suite. The current quantitative perfusion imaging includes two cascaded steps, time-resolved image reconstruction and perfusion parametric estimation. For time-resolved image reconstruction, technical challenge imposed by poor temporal resolution and sampling density causes inaccurate quantification of the temporal variation of cerebral artery and tissue attenuation values. For perfusion parametric estimation, it remains a technical challenge to appropriately handcraft the regularization for better solving the perfusion parametric estimation problem. These two challenges together prevent obtaining quantitatively accurate perfusion images using C-arm CBCT. The purpose of this work is to simultaneously address these two challenges via combining the two cascaded steps into a single joint optimization problem and reconstructing quantitative perfusion images directly from the measured sinogram data. In the developed direct cerebral perfusion parametric image reconstruction technique, TRAINER in short, the quantitative perfusion images have been represented as a subject-specific conditional generative model trained under the constraint of the time-resolved CT forward model, perfusion convolutional model, and the subject’s own measured sinogram data. Results shown in this paper demonstrated that using TRAINER, quantitative cerebral perfusion images can be accurately obtained using C-arm CBCT in the interventional suite.
@@ I-B Technical Challenges for Quantitative Cerebral Perfusion Imaging Using C-arm Cone-Beam CT
Several techniques have previously been proposed to address the potential challenges in time-resolved image reconstruction using C-arm CBCT. One scheme [7, 8, 9] was to decompose the time attenuation curves (TACs) of different image voxels in terms of known temporal basis functions such as Gaussian functions, gamma-variate functions, linear or cubic splines. The incorporation of known basis functions dramatically reduces the number of unknowns, hence, strongly regularizes the otherwise ill-posed optimization problem. In the second scheme, a temporal deconvolution method was proposed in [10] to improve the temporal resolution and temporal sampling density by deconvolving the reconstructed TACs with a known convolution kernel simulating the temporal resolution degradation scenario. In the third scheme, a local TACs optimization framework was proposed in [11] to model the TACs inside the vessel regions and optimize the TACs constrained by the acquired data. The fourth scheme was to improve temporal resolution via the limited-view image reconstruction strategy. One such method is the SMART-RECON [12], which reconstructs up to four time frames with temporal variation without suffering from severe limited-view artifacts using data acquired over the short-scan range [13]. By incorporating the prior knowledge of periodicity of the limited-view artifacts in the C-arm CBCT perfusion data acquisition scheme, eSMART-RECON [6] further improves the temporal resolution to 4-7.5 frames per second (fps). Additionally, by recasting the temporal resolution improvement task as a temporal extrapolation task, AIRPORT [14] improves the temporal resolution to 40 fps, the time window needed to acquire a single projection view data using typical C-arm CBCT and the data acquired over the short-scan range.
@@ I-D Technical Challenges for Perfusion Parametric Estimation
Perfusion parametric estimation aims to quantify cerebral blood flow (CBF), cerebral blood volume (CBV), mean transit time (MTT), and time to peak (TTP) etc using the reconstructed time-resolved CT images. For perfusion parametric estimation, currently available techniques [15, 16, 17, 18, 19, 20, 21] incorporate handcrafted regularization to heavily constrain the otherwise ill-posed deconvolution problem especially under non-ideal data acquisition conditions such as low exposure levels. However, it remains a technical challenge to appropriately design the regularization for better solving the deconvolution problem. Inappropriate designs of regularizations may lead to bias in the obtained parametric images [22, 23], even if the used time-resolved images are quantitatively accurate.
@@ I-E Purpose and Innovations of This Work
Although significant efforts have been made to reconstruct time-resolved images and estimate perfusion parametric images from the reconstructed time-resolved CT images, no work has yet been published to address inaccurate perfusion imaging jointly caused by the poor temporal resolution of C-arm CBCT and current perfusion parametric estimation techniques due to biased regularization. These two challenges together prevent the acquisition of quantitatively accurate perfusion images using C-arm CBCT.
In the current quantitative perfusion imaging pipeline, time-resolved image reconstruction and perfusion parametric estimation are treated as two cascaded challenges. It is our conviction that addressing each of them independently may fail to achieve the needed overall accuracy. On the contrary, in this work, instead of addressing each challenge independently, we recast the quantitative perfusion imaging problem as a single joint optimization problem and reconstruct quantitative perfusion images directly from the measured sinogram data of each individual subject. The proposed single joint optimization strategy has two unique advantages. First, it significantly reduces the number of unknowns to be determined from the number of time frames of the time-resolved images to the number of perfusion parametric images, resulting in a powerful regularization to the otherwise ill-posed optimization problems. Second, the upper bound of the quantitative accuracy of perfusion images reconstructed via the proposed single joint optimization strategy only depends on the time window needed to acquire a single projection view data, not the temporal resolution of tomographic reconstruction (a.k.a., the gantry rotation speed), relaxing the requirements of C-arm rotation speed, mechanical stability or other system conditions.
To implement this idea, we developed direct cerebral perfusion parametric image reconstruction technique, TRAINER in short, that jointly leverages the forward model of the time-resolved CT data acquisition and the convolutional model for cerebral perfusion parametric estimation. To avoid potentially biased regularization design for perfusion parametric estimation, the perfusion parametric images in TRAINER have been represented as a subject-specific conditional generative model trained under the constraint of the subject’s own measured sinogram data.
@@
  • It explicitly leverages the measured sinogram data of each individual subject to ensure the accuracy of perfusion parametric images for each individual subject;
@@
In this work, we demonstrated that using TRAINER, quantitative perfusion images can be accurately obtained via training a subject-specific conditional generative model under the constraint of the subject’s own measured sinogram data. Results shown in the paper demonstrated that the two technical challenges, i.e., poor temporal resolution of C-arm CBCT and inaccurate perfusion parametric estimation methods, have been simultaneously addressed using the proposed TRAINER technique. TRAINER may help quickly identify patients who are most likely to benefit from EVT directly using C-arm CBCT in the interventional suite.
@@ II-A Forward Model for Time-Resolved CT Data Acquisition
Considering the ideal CT imaging scenario where the image object is static and stationary and a monochromatic X-ray spectrum is used in the data acquisition, the forward model under proper digitization can be formulated as the following linear system:
$\bar{\mathbf{y}}$ $=\mathbf{A}\mathbf{x}$ (1)
@@
$\bar{\mathbf{y}}_{t}$ $=\mathbf{A}_{t}\mathbf{x}_{t}$ (2)
Where $\bar{\mathbf{y}}_{t}\in\mathbb{R}^{\frac{M}{K}\times 1}$, $\mathbf{x}_{t}\in\mathbb{R}^{N\times 1}$ denote the modeled line integral values and the image to be reconstructed associated with the $t$-th time frame respectively, $\mathbf{A}_{t}\in\mathbb{R}^{\frac{M}{K}\times N}$ denotes CT system matrix corresponding to $t$-th time frame, where $t=1,2,\cdots,T$ denotes the index of time frame. Here $T_{0}$ denotes the number of short-scan acquisitions, $K$ denotes the number of subsets within a single short-scan range, and $T=KT_{0}$ denotes the total number of time frames to be reconstructed.
Since our goal is to jointly reconstruct the dynamic image object $\{\mathbf{x}_{t}\}$, it is more convenient to stack these linear sub-systems as follows:
@@
$\mathbf{X}$ $=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{T}]$ (5)
$\bar{\mathbf{Y}}$ $=[\bar{\mathbf{y}}_{1},\bar{\mathbf{y}}_{2},\cdots,\bar{\mathbf{y}}_{T}]$ (6)
where $\mathcal{A}\in\mathbb{R}^{MT_{0}\times NT}$ takes each system matrix $\mathbf{A}_{t}$ as its diagonal matrix block, $\mathbf{X}\in\mathbb{R}^{N\times T}$ denotes the time-resolved image matrix stacking each image associated with the $t$-th time frame $\mathbf{x}_{t}$ along its column dimension, and $\vec{\cdot}$ denotes the vectorization operator.
@@ II-B Convolutional Model for Cerebral Perfusion Parametric Estimation
In patients with an AIS, the TACs may lag behind the AIF, particularly if the latter is obtained from unaffected arteries in the contralateral hemisphere [24]. To account for this delay, a single-compartment model is utilized to incorporate the lag effect of tissue TACs, expressing the impulse residue function as an exponential form associated with the perfusion parameter CBF and $\mathbf{\mathcal{T}}_{0}$ (time lag). $\mathbf{\mathcal{T}}_{0}$ serves as an additional valuable parameter for characterizing hemodynamic status in cerebral circulation of AIS patients [24]. According to the convolutional model for cerebral perfusion parametric mapping [24, 25, 26, 27], the relationship between the time-resolved image matrix $\mathbf{X}$ and the flow-scaled residual matrix $\mathbf{C}$ can be represented by the following linear model:
$\mathbf{X}$ $=\Delta t\cdot\mathbf{C}\mathbf{B}^{\text{T}}$ (7)
$B_{i,j}$ =$=\begin{cases}h_{\text{AIF}}(t_{i-j+1}),&j\leq i\\[5.0pt] 0,&j>i\end{cases}$ (8)
$C_{i,j}$ =$=\begin{cases}\rho\cdot\text{CBF}_{i},&t_{j}<\mathcal{T}_{0_{i}}\\[2.0pt] \rho\cdot\text{CBF}_{i}\cdot\exp\left(-\frac{\left(t_{j}-\mathcal{T}_{0_{i}}\right)}{h}\right),&t_{j}\geq\mathcal{T}_{0_{i}}\end{cases}$ (9)
Where $\mathbf{B}\in\mathbb{R}^{T\times T}$, $B_{i,j}$ denotes the matrix element of $\mathbf{B}$, $h_{\text{AIF}}(t_{i})$ denotes the digitized arterial input function [28]. $\Delta t$ denotes the unit time interval, and $h$ denotes a unit conversion constant with a unit of seconds. $C_{i,j}$ denotes the matrix element of $\mathbf{C}\in\mathbb{R}^{N\times T}$ which represents the flow-scaled residual matrix stacking up each flow-scaled residual function values of all image voxels associated with the $j$-th time frame, $\mathbf{c}_{j}\in\mathbb{R}^{N\times 1}$, along its column direction. The matrix form $\mathbf{C}\mathbf{B}^{\text{T}}$ represents the digitized convolutional process of the arterial input function $h_{\text{AIF}}(t)$ and the flow-scaled residual function for each image voxel independently. $\rho$ denotes the nominal density of cerebral tissues. $\text{CBF}_{i}$ and $\mathcal{T}_{0_{i}}$ denote the $i$-th element of the two perfusion parametric images to be reconstructed, CBF and $\mathbf{\mathcal{T}}_{0}$, respectively.
According to the convolutional model for cerebral perfusion parametric mapping, the perfusion parametric images CBV and MTT can be derived using the following equations once the optimal CBF and $\mathbf{\mathcal{T}}_{0}^{*}$ are obtained:
$\text{CBV}^{*}_{i}$ $\approx\frac{1}{\rho}\cdot\sum_{j=1}^{T}[\mathbf{C}(\text{CBF}^{*},\mathbf{\mathcal{T}}_{0}^{*})]_{i,j}\cdot\Delta t$ (10)
$\text{MTT}^{*}_{i}$ $=\frac{\text{CBV}^{*}_{i}}{\text{CBF}^{*}_{i}}$ (11)
$\mathbf{\text{CBV}}^{*}_{i}$ and $\mathbf{\text{MTT}}^{*}_{i}$ denote the $i$-th element of the two obtained perfusion parametric images, CBV and MTT, respectively.
@@ II-C Quantitative Cerebral Perfusion Image Reconstruction Directly From Measured Sinogram Data
The two perfusion parametric images to be reconstructed, CBF and $\mathbf{\mathcal{T}}_{0}$, can be obtained by solving the following optimization problem:
@@
$\mathbf{U}$ $:=\left[\mathbf{\text{CBF}},\mathbf{\mathcal{T}}_{0}\right],$
where $\mathbf{W}\in\mathbb{R}^{MT_{0}\times MT_{0}}$ denotes the statistical weighting matrix and $\vec{\mathbf{Y}}\in\mathbb{R}^{MT_{0}\times 1}$ denotes the measured data stacking the measurements associated with each scan along its column direction.
@@ II-D Subject-Specific Conditional Generative Models for Perfusion Parametric Image Representation
To avoid potentially biased regularization design for perfusion parametric estimation, the perfusion parametric images in the proposed TRAINER have been represented as a subject-specific conditional generative model trained under the constraint of the subject’s own measured sinogram data. We parameterize the unknowns (CBF and $\mathbf{\mathcal{T}}_{0}$) as follows:
@@
where $\mathcal{G}_{\mathbf{\Theta}}$ denotes the parameterized perfusion images, CBF and $\mathbf{\mathcal{T}}_{0}$, from the noise image sampled from the uniform distribution, $\mathbf{Z}\in\mathbb{R}^{N\times 2}$. Inspired by the deep image prior [29], we represent $\mathcal{G}_{\mathbf{\Theta}}$ by typical convolutional neural networks, whose model parameters are denoted by $\mathbf{\Theta}$.